# Tree-Guided Diffusion Planner

### v1

@@ Abstract
Planning with pretrained diffusion models has emerged as a promising approach for solving test-time guided control problems. However, standard gradient guidance typically performs optimally under convex and differentiable reward landscapes, showing substantially reduced effectiveness in real-world scenarios involving non-convex objectives, non-differentiable constraints, and multi-reward structures. Furthermore, recent supervised planning approaches require task-specific training or value estimators, which limits test-time flexibility and zero-shot generalization. We propose a Tree-guided Diffusion Planner (TDP), a zero-shot test-time planning framework that balances exploration and exploitation through structured trajectory generation. We frame test-time planning as a tree search problem using a bi-level sampling process: (1) diverse parent trajectories are produced via training-free particle guidance to encourage broad exploration, and (2) sub-trajectories are refined through fast conditional denoising guided by task objectives. TDP addresses the limitations of gradient guidance by exploring diverse trajectory regions and harnessing gradient information across this expanded solution space using only pretrained models and test-time reward signals. We evaluate TDP on three diverse tasks: maze gold-picking, robot arm block manipulation, and AntMaze multi-goal exploration. TDP consistently outperforms state-of-the-art approaches on all tasks. The project page can be found at: tree-diffusion-planner.github.io.
@@ 1 Introduction
Diffusion models offer a data-driven framework for planning, enabling the generation of coherent and expressive trajectories learning from offline demonstrations Janner et al. (2022); Liang et al. (2023); Li et al. (2023); Chen et al. (2024b). Compared to single-step model-free reinforcement learning (RL) methods Kumar et al. (2020); Kostrikov et al. (2021), diffusion planners are more effective for long-horizon planning by generating temporally extended trajectories through multi-step prediction. Without task-specific dynamics models, pretrained diffusion planners can be adapted for test-time planning through guidance functions that provide numerical scores for user requirements such as state conditions or physical constraints. Prior studies in diffusion guidance algorithm (e.g., classifier guidance Dhariwal and Nichol (2021)) have been successfully incorporated to generate conditional trajectory samples given test-time reward signals Janner et al. (2022); Ajay et al. (2023). These guidance algorithms present an exploration-exploitation trade-off Sutton and Barto (1998), balancing adherence to the pretrained model for feasibility against maximizing the external guide score that may require exploring out-of-distribution trajectories.
@@
As the main bottleneck in guided planning lies in the limited quality of trajectory samples produced by the pretrained models, prior works have primarily focused on improving general sample quality and mitigating planning artifacts, while the guidance algorithms themselves remain relatively underexplored. The majority of recent studies on diffusion planners emphasize advancing supervised planning capabilities Chen et al. (2024a); Yoon et al. (2025); Li et al. (2023); Chen et al. (2024b); Liang et al. (2023); Zhou et al. (2025). They are categorized into sequential Chen et al. (2024a); Yoon et al. (2025), hierarchical Li et al. (2023); Chen et al. (2024b), and fine-tuning Liang et al. (2023); Zhou et al. (2025) approaches. These works improve the modeling of the underlying system dynamics from the static offline RL benchmarks. Recent works successfully solve challenging offline benchmarks by reformulating the training scheme, learning value estimator, and test-time scaling. Another recent research direction in diffusion planners aims to enhance zero-shot planning capabilities Wang et al. (2025). Test-time tasks are shifted from the training distribution, and planners are given access to a pretrained model along with dense reward signals to adapt to these unseen tasks effectively.
@@
However, test-time planning capabilities are often evaluated in relatively simple optimization tasks, such as minimizing the distance to optimal trajectories or matching the outputs of pretrained classifiers trained on expert demonstrations Janner et al. (2022); Chen et al. (2024b), and predominantly within in-distribution trajectory settings Chen et al. (2024b); Liang et al. (2023). Typical benchmarks include maze navigation tasks Chen et al. (2024a) where agents minimize distance to a single goal, or block stacking tasks Janner et al. (2022) where a unique target configuration maximizes reward. These tasks generally involve convex optimization problems, where a unique global optimal trajectory maximizes a smooth guide function. In addition, similar to model-free RL methods Wang et al. (2020), guided planning often relies on a pretrained value estimator trained on supervised trajectory data; however, collecting optimal trajectories for each new task is often impractical or infeasible, particularly for tasks with complex dynamics or limited simulation control.
@@
In this work, we address a fundamental challenge of existing diffusion-based planners: while they excel at generating low-level action sequences, most real-world planning tasks require decision-making over high-level abstractions. These abstractions often introduce non-convex guide functions or non-differentiable constraints, making them incompatible with conventional test-time guidance methods that assume smooth, convex optimization landscapes. For example, maze navigation with intermediate goal bypassing Liang et al. (2023) introduces a non-differentiable rule into the planning process. In multi-reward block stacking, the agent must reconcile multiple configuration-dependent reward signals to determine the most favorable block arrangement. These scenarios underscore the need for guided planning algorithms that flexibly accommodate complex test-time specifications while ensuring both trajectory feasibility and task-specific fitness. Recently, several training-free guidance algorithms have been proposed in the image domain Chung et al. (2023); Song et al. (2023); Yu et al. (2023); Guo et al. (2024), but their capabilities are typically restricted to smooth differentiable guide functions. Since test-time guide functions can take any form, there is a need for a flexible planning algorithm capable of handling a broad class of guide functions.
@@
We propose Tree-guided Diffusion Planner (TDP), which formulates test-time planning as a tree search problem balancing exploration through diverse trajectory samples and exploitation via guided sub-trajectories. While pretrained diffusion planners model underlying system dynamics, TDP samples high-reward (i.e., high guidance score for the test task) solution trajectories conditioned on the learned dynamics in a zero-shot manner. TDP equips the pretrained diffusion planner model with the ability to reason over higher-level objectives. TDP outperforms state-of-the-art planning methods on challenging tasks with non-convex guide functions and non-differentiable constraints across all test scenarios. TDP enables flexible task-aware planning without requiring expert demonstrations.
@@ 2 Related Work
  • Sequential approaches Chen et al. (2024a); Yoon et al. (2025) explore one action at a time, which works well in single-goal convex tasks but struggles with multi-goal tasks where different goals have varying test-time priorities. In such settings, they often converge to local optima instead of discovering distant, high-priority goals. Furthermore, sequential approaches like MCTD Yoon et al. (2025) typically require training task-specific value estimators to guide their single-step decisions, limiting applicability to zero-shot scenarios where no task-specific training is available. In contrast, TDP performs multi-step exploration through diverse bi-level trajectory sampling, which better handles challenging multi-goal scenarios without requiring additional training components beyond the pretrained planner.
  • Hierarchical diffusion planners Li et al. (2023); Chen et al. (2024b) rely on training-time supervision to learn sub-goal distributions. They perform well when both initial and goal states are given, but struggle on unlabeled zero-shot tasks such as the test-time gold-picking task Liang et al. (2023). In standard maze navigation benchmarks, Hierarchical Diffuser Chen et al. (2024b) tends to generate shortest-path trajectories when initial and goal states are specified. However, the gold-picking task poses a fundamentally different challenge: the goal is hidden and often misaligned with the shortest path.
  • Diffusion model predictive control (D-MPC) Zhou et al. (2025) adapts to changing dynamics via few-shot fine-tuning with expert demonstrations, but struggles with unseen long-horizon tasks and complex behaviors as standard dynamics models $p(s|a)$ struggle to capture long-context reward structures effectively. In contrast, TDP models the joint distribution $p(s,a)$ to enable solving long-horizon and multi-goal tasks, and is a fully zero-shot planner, operating without test-time demonstrations.
@@
Training-free guidance methods leverage structural priors and domain knowledge for control without additional learning. Classical planners (e.g., A*, potential fields) Hart et al. (1968); Dijkstra (1959); Tang et al. (2010); Geraerts and Overmars (2004); Cortés and Siméon (2005) compute feasible trajectories through graph search or geometric reasoning. In continuous domains, trajectory optimization and model predictive control Amos et al. (2019); Schwenzer et al. (2021) refine actions iteratively using known dynamics. Local search techniques (e.g., hill climbing), guided policy search, and reward shaping Levine and Koltun (2013); Trott et al. (2019); Hu et al. (2020) serve as strong baselines for structured, domain-specific control problems but typically lack the flexibility to generalize beyond narrow task settings. In contrast, diffusion-based test-time planning targets complex tasks that need out-of-distribution generalization and adaptability.
@@
Tree structures naturally represent hierarchical sequential decisions. Trajectory Aggregation Tree (TAT) Feng et al. (2024) mitigates artifacts in diffusion-generated trajectories by aggregating similar states near the initial state, but it struggles with complex long-horizon dependencies due to its limited aggregation depth early in the trajectory. Monte Carlo Tree Search Świechowski et al. (2022); Hennes and Izzo (2015); Labbé et al. (2020); Guo et al. (2014) explores full-horizon trajectories via stochastic roll-outs, but its reliance on discrete actions and reward heuristics limits scalability in high-dimensional or continuous control tasks with external guidance. In contrast, TDP’s bi-level tree framework integrates gradient-based guidance at both parent and child levels, enabling structured and adaptive planning under complex test-time objectives.
@@ 3.1 Problem Setting
We consider the test-time reward maximization problem on a discrete-time dynamics system $\boldsymbol{s}_{t+1}=f(\boldsymbol{s}_{t},\boldsymbol{a}_{t})$ via pretrained planner model, where the agent has access to the user-defined guide function $\mathcal{J}(\boldsymbol{\tau})$, which indicates the fitness of the generated trajectory $\boldsymbol{\tau}$. As per-timestep reward does not guarantee the optimality of a low-level action (e.g., non-convex reward landscape), planning capability based on exploration is required to find the optimal trajectory $\hat{\boldsymbol{\tau}}$ that maximizes $\mathcal{J}$. The agent must find an action sequence that maximizes the guide score within a limited number of steps:
@@
Planning horizon $T_{\text{pred}}$ is determined by the choice of planner model. Model-free RL methods with single step execution Kumar et al. (2020); Kostrikov et al. (2021) predict a single action at each timestep so $T_{\text{pred}}=1$, whereas diffusion planner Janner et al. (2022) predicts a sequence of actions $\boldsymbol{a}_{1:T_{\text{pred}}}$ at once. As diffusion planners predict more future states, they benefit from capturing longer-term contextual information. For example, in a long-horizon multi-goal navigation task Pitis et al. (2020), an agent may encounter several intermediate suboptimal goals, but reaching a farther goal yields a substantially larger reward, requiring planning several steps ahead rather than greedy pursuit of nearby rewards. Planning over longer horizons can enhance performance on more challenging tasks, particularly when future rewards are more significant Kaiser et al. (2024); Schrittwieser et al. (2020).
@@ 3.2 Test-time Guided Planning with Diffusion Models
The standard approach to guide diffusion planning in test time is to use naïve gradient guidance Guo et al. (2024), which progressively refines the denoising process by combining the score estimate from the unconditional diffusion model with the auxiliary guide function Janner et al. (2022); Chen et al. (2024b). It approximates the reverse denoising process as Gaussian with small perturbation if the guidance distribution $h(\boldsymbol{\tau}_{i})$ is sufficiently smooth and the gradient of the guide function is time-independent:
@@
where $g=\nabla_{\tau}\log h(\boldsymbol{\tau}_{i})$ is the gradient of the guidance distribution Sohl-Dickstein et al. (2015), $\alpha$ is guidance strength, and $\mu,\Sigma$ are the mean and covariance of the pretrained reverse denoising process. On the other hand, classifier guidance (CG) Dhariwal and Nichol (2021) and classifier-free guidance (CFG) Ho and Salimans (2022) are also broadly used in guided planning Ajay et al. (2023); Li et al. (2023); Zhou et al. (2023). However, they require access to expert demonstration data to train either time-dependent classifier model or conditional diffusion planner model based on trajectory rewards for given task. Despite their simple yet powerful architecture, it is expensive to extend CG and CFG in test-time as it requires collecting expert demonstration for each new task and retraining the model. The ultimate goal of test-time guided planning is adaptive planning, identifying trajectories that satisfy user requirements using a pretrained diffusion planner without additional expert supervision.

### v2

@@ Abstract
Planning with pretrained diffusion models has emerged as a promising approach for solving test-time guided control problems. Standard gradient guidance typically performs optimally under convex, differentiable reward landscapes. However, it shows substantially reduced effectiveness in real-world scenarios with non-convex objectives, non-differentiable constraints, and multi-reward structures. Furthermore, recent supervised planning approaches require task-specific training or value estimators, which limits test-time flexibility and zero-shot generalization. We propose a Tree-guided Diffusion Planner (TDP), a zero-shot test-time planning framework that balances exploration and exploitation through structured trajectory generation. We frame test-time planning as a tree search problem using a bi-level sampling process: (1) diverse parent trajectories are produced via training-free particle guidance to encourage broad exploration, and (2) sub-trajectories are refined through fast conditional denoising guided by task objectives. TDP addresses the limitations of gradient guidance by exploring diverse trajectory regions and harnessing gradient information across this expanded solution space using only pretrained models and test-time reward signals. We evaluate TDP on three diverse tasks: maze gold-picking, robot arm block manipulation, and AntMaze multi-goal exploration. TDP consistently outperforms state-of-the-art approaches on all tasks. The project page can be found at: tree-diffusion-planner.github.io.
@@ 1 Introduction
Diffusion models offer a data-driven framework for planning, enabling the generation of coherent and expressive trajectories learning from offline demonstrations [janner2022planningdiffusionflexiblebehavior, liang2023adaptdiffuserdiffusionmodelsadaptive, pmlrv202li23ad, chen2024simplehierarchicalplanningdiffusion]. Compared to single-step model-free reinforcement learning (RL) methods [kumar2020conservativeqlearningofflinereinforcement, kostrikov2021offlinereinforcementlearningimplicit], diffusion planners are more effective for long-horizon planning by generating temporally extended trajectories through multi-step prediction. Without task-specific dynamics models, pretrained diffusion planners can be adapted for test-time planning through guidance functions that provide numerical scores for user requirements such as state conditions or physical constraints. Prior studies in diffusion guidance algorithm (e.g., classifier guidance [dhariwal2021diffusionmodelsbeatgans]) have been successfully incorporated to generate conditional trajectory samples given test-time reward signals [janner2022planningdiffusionflexiblebehavior, ajay2023conditionalgenerativemodelingneed]. These guidance algorithms present an exploration-exploitation trade-off [712192], balancing adherence to the pretrained model for feasibility against maximizing the external guide score, which may require exploring out-of-distribution trajectories.
@@
As the main bottleneck in guided planning lies in the limited quality of trajectory samples produced by the pretrained models, prior works have primarily focused on improving general sample quality and mitigating planning artifacts, while the guidance algorithms themselves remain relatively underexplored. The majority of recent studies on diffusion planners emphasize advancing supervised planning capabilities [liang2023adaptdiffuserdiffusionmodelsadaptive, pmlrv202li23ad, chen2024simplehierarchicalplanningdiffusion, chen2024diffusionforcingnexttokenprediction, yoon2025montecarlotreediffusion, zhou2025diffusionmodelpredictivecontrol]. They are categorized into sequential [chen2024diffusionforcingnexttokenprediction, yoon2025montecarlotreediffusion], hierarchical [pmlrv202li23ad, chen2024simplehierarchicalplanningdiffusion], and fine-tuning [liang2023adaptdiffuserdiffusionmodelsadaptive, zhou2025diffusionmodelpredictivecontrol] approaches. These works improve the modeling of the underlying system dynamics from the static offline RL benchmarks. Recent works successfully solve challenging offline benchmarks by reformulating the training scheme, learning value estimator, and test-time scaling. Another recent research direction in diffusion planners aims to enhance zero-shot planning capabilities [wang2025inferencetimepolicysteeringhuman]. Test-time tasks are shifted from the training distribution, and planners are given access to a pretrained model along with dense reward signals to adapt to these unseen tasks effectively.
@@
However, test-time planning capabilities are often evaluated in relatively simple optimization tasks, such as minimizing the distance to optimal trajectories or matching the outputs of pretrained classifiers trained on expert demonstrations [janner2022planningdiffusionflexiblebehavior, chen2024simplehierarchicalplanningdiffusion], and predominantly within in-distribution trajectory settings [liang2023adaptdiffuserdiffusionmodelsadaptive, chen2024simplehierarchicalplanningdiffusion]. Typical benchmarks include maze navigation tasks [chen2024diffusionforcingnexttokenprediction] where agents minimize distance to a single goal, or block stacking tasks [janner2022planningdiffusionflexiblebehavior] where a unique target configuration maximizes reward. These tasks generally involve convex optimization problems, where a unique global optimal trajectory maximizes a smooth guide function. In addition, similar to model-free RL methods [wang2020benchmarking], guided planning often relies on a pretrained value estimator trained on supervised trajectory data; however, collecting optimal trajectories for each new task is often impractical or infeasible, particularly for tasks with complex dynamics or limited simulation control.
@@
In this work, we address a fundamental challenge of existing diffusion-based planners: while they excel at generating low-level action sequences, most real-world planning tasks require decision-making over high-level abstractions. These abstractions often introduce non-convex guide functions or non-differentiable constraints, making them incompatible with conventional test-time guidance methods that assume smooth, convex optimization landscapes. For example, maze navigation with intermediate goal bypassing [liang2023adaptdiffuserdiffusionmodelsadaptive] introduces a non-differentiable rule into the planning process. In multi-reward block stacking, the agent must reconcile multiple configuration-dependent reward signals to determine the most favorable block arrangement. These scenarios underscore the need for guided planning algorithms that flexibly accommodate complex test-time specifications while ensuring both trajectory feasibility and task-specific fitness. Recently, several training-free guidance algorithms have been proposed in the image domain [chung2023diffusion, pmlrv202song23k, yu2023freedomtrainingfreeenergyguidedconditional, guo2024gradientguidancediffusionmodels], but their capabilities are typically restricted to smooth differentiable guide functions. Since test-time guide functions can take any form, there is a need for a flexible planning algorithm that can handle a broad class of guide functions.
@@
We propose Tree-guided Diffusion Planner (TDP), which formulates test-time planning as a tree search problem that balances exploration via diverse trajectory samples and exploitation via guided sub-trajectories. While pretrained diffusion planners model underlying system dynamics, TDP samples high-reward (i.e., high guidance score for the test task) solution trajectories conditioned on the learned dynamics in a zero-shot manner. TDP equips the pretrained diffusion planner model with the ability to reason over higher-level objectives. TDP outperforms state-of-the-art planning methods on challenging tasks with non-convex guide functions and non-differentiable constraints across all test scenarios. TDP enables flexible task-aware planning without requiring expert demonstrations.
@@ 2 Related Work
  • Sequential approaches [chen2024diffusionforcingnexttokenprediction, yoon2025montecarlotreediffusion] explore one action at a time, which works well in single-goal convex tasks but struggles with multi-goal tasks where different goals have varying test-time priorities. In such settings, they often converge on local optima rather than discovering distant, high-priority goals. Furthermore, sequential approaches like MCTD [yoon2025montecarlotreediffusion] typically require training task-specific value estimators to guide their single-step decisions, limiting applicability to zero-shot scenarios where no task-specific training is available. In contrast, TDP performs multi-step exploration through diverse bi-level trajectory sampling, which better handles challenging multi-goal scenarios without requiring additional training components beyond the pretrained planner.
  • Hierarchical diffusion planners [pmlrv202li23ad, chen2024simplehierarchicalplanningdiffusion] rely on training-time supervision to learn sub-goal distributions. They perform well when both initial and goal states are given, but struggle on unlabeled zero-shot tasks such as the test-time gold-picking task [liang2023adaptdiffuserdiffusionmodelsadaptive]. In standard maze navigation benchmarks, Hierarchical Diffuser [chen2024simplehierarchicalplanningdiffusion] tends to generate shortest-path trajectories when initial and goal states are specified. However, the gold-picking task poses a fundamentally different challenge: the goal is hidden and often misaligned with the shortest path.
  • Diffusion model predictive control (D-MPC) [zhou2025diffusionmodelpredictivecontrol] adapts to changing dynamics via few-shot fine-tuning with expert demonstrations, but struggles with unseen long-horizon tasks and complex behaviors as standard dynamics models $p(s|a)$ struggle to capture long-context reward structures effectively. In contrast, TDP models the joint distribution $p(s,a)$ to enable solving long-horizon and multi-goal tasks and is a fully zero-shot planner that operates without test-time demonstrations.
@@
Training-free guidance methods leverage structural priors and domain knowledge for control without additional learning. Classical planners (e.g., A*, potential fields) [4082128, dijkstra1959note, 5565069, Geraerts2004, Cortés2005] compute feasible trajectories through graph search or geometric reasoning. In continuous domains, trajectory optimization and model predictive control [amos2019differentiablempcendtoendplanning, Schwenzer_Ay_Bergs_Abel_2021] refine actions iteratively using known dynamics. Local search techniques (e.g., hill climbing), guided policy search, and reward shaping [pmlr-v28-levine13, trott2019keepingdistancesolvingsparse, hu2020learningutilizeshapingrewards] serve as strong baselines for structured, domain-specific control problems but typically lack the flexibility to generalize beyond narrow task settings. In contrast, diffusion-based test-time planning targets complex tasks that need out-of-distribution generalization and adaptability.
@@
Tree structures naturally represent hierarchical sequential decisions. Trajectory Aggregation Tree (TAT) [feng2024resistingstochasticrisksdiffusion] mitigates artifacts in diffusion-generated trajectories by aggregating similar states near the initial state, but it struggles with complex long-horizon dependencies due to its limited aggregation depth early in the trajectory. Monte Carlo Tree Search [_wiechowski_2022, hennes2015interplanetary, 9036915, NIPS2014_88bf0c64] explores full-horizon trajectories via stochastic roll-outs, but its reliance on discrete actions and reward heuristics limits scalability in high-dimensional or continuous control tasks with external guidance. In contrast, TDP’s bi-level tree framework integrates gradient-based guidance at both parent and child levels, enabling structured and adaptive planning under complex test-time objectives.
@@ 3.1 Problem Setting
We consider the test-time reward maximization problem on a discrete-time dynamics system $\boldsymbol{s}_{t+1}=f(\boldsymbol{s}_{t},\boldsymbol{a}_{t})$ via a pretrained planner model, where the agent has access to the user-defined guide function $\mathcal{J}(\boldsymbol{\tau})$, which indicates the fitness of the generated trajectory $\boldsymbol{\tau}$. As per-timestep reward does not guarantee the optimality of a low-level action (e.g., non-convex reward landscape), planning capability based on exploration is required to find the optimal trajectory $\hat{\boldsymbol{\tau}}$ that maximizes $\mathcal{J}$. The agent must find an action sequence that maximizes the guide score within a limited number of steps:
@@
Planning horizon $T_{\text{pred}}$ is determined by the choice of planner model. Model-free RL methods with single step execution [kumar2020conservativeqlearningofflinereinforcement, kostrikov2021offlinereinforcementlearningimplicit] predict a single action at each timestep so $T_{\text{pred}}=1$, whereas diffusion planner [janner2022planningdiffusionflexiblebehavior] predicts a sequence of actions $\boldsymbol{a}_{1:T_{\text{pred}}}$ at once. As diffusion planners predict more future states, they benefit from capturing longer-term contextual information. For example, in a long-horizon multi-goal navigation task [pitis2020maximumentropygainexploration], an agent may encounter several intermediate suboptimal goals, but reaching a farther goal yields a substantially larger reward, requiring planning several steps rather than greedy pursuit of nearby rewards. Planning over longer horizons can enhance performance on more challenging tasks, particularly when future rewards are more significant [kaiser2024modelbasedreinforcementlearningatari, Schrittwieser_2020].
@@ 3.2 Test-time Guided Planning with Diffusion Models
The standard approach to guide diffusion planning in test time is to use naïve gradient guidance [guo2024gradientguidancediffusionmodels], which progressively refines the denoising process by combining the score estimate from the unconditional diffusion model with the auxiliary guide function [janner2022planningdiffusionflexiblebehavior, chen2024simplehierarchicalplanningdiffusion]. It approximates the reverse denoising process as Gaussian with small perturbation if the guidance distribution $h(\boldsymbol{\tau}_{i})$ is sufficiently smooth and the gradient of the guide function is time-independent:
@@
where $g=\nabla_{\tau}\log h(\boldsymbol{\tau}_{i})$ is the gradient of the guidance distribution [sohldickstein2015deepunsupervisedlearningusing], $\alpha$ is guidance strength, and $\mu,\Sigma$ are the mean and covariance of the pretrained reverse denoising process. On the other hand, classifier guidance (CG) [dhariwal2021diffusionmodelsbeatgans] and classifier-free guidance (CFG) [ho2022classifierfreediffusionguidance] are also broadly used in guided planning [pmlrv202li23ad, ajay2023conditionalgenerativemodelingneed, zhou2023adaptive]. However, they require access to expert demonstration data to train either a time-dependent classifier model or a conditional diffusion planner model based on trajectory rewards for a given task. Despite their simple yet powerful architecture, it is expensive to extend CG and CFG in test-time as it requires collecting expert demonstrations for each new task and retraining the model. The ultimate goal of test-time guided planning is adaptive planning, identifying trajectories that satisfy user requirements using a pretrained diffusion planner without additional expert supervision.